# Predictive Data Selection: The Data That Predicts Is the Data That Teaches

### v3

@@ Abstract
Language model pretraining involves training on extensive corpora, where data quality plays a pivotal role. In this work, we aim to directly estimate the contribution of data during pretraining and select pretraining data in an efficient manner. Specifically, we draw inspiration from recent findings showing that compression efficiency (i.e., the normalized loss) of diverse models on certain text correlates strongly with their downstream performance, when the text domain aligns with the downstream benchmarks (Huang et al., 2024). Building on this observation, we hypothesize that data on which model losses are predictive of downstream abilities also contribute effectively to learning. To leverage this insight, we introduce predictive data selection (PreSelect), a lightweight and efficient data selection method that requires training and deploying only a fastText-based scorer. Through comprehensive experiments with 1B and 3B parameter models, we demonstrate that models trained on 30B tokens selected with PreSelect surpass the performance of the vanilla baseline trained on 300B tokens, achieving a 10x reduction in compute requirements. Furthermore, PreSelect significantly outperforms other competitive data selection baselines, such as DCLM and FineWeb-Edu on a scale of 3B models trained on 100B tokens. We open-source our trained data selection scorer along with the curated datasets at https://github.com/hkust-nlp/PreSelect.
@@ 1 Introduction
Large language model (LLM) pre-training typically requires training on a huge data source such as web crawl data where the existence of low-quality data leads to slow scaling law (Kaplan et al., 2020; Chowdhery et al., 2023). Given the growing available token budget and limited training budget nowadays, data selection has become a standard step in the pre-training stage. This step has played a critical role in enhancing data quality and optimizing training efficiency during the development of various LLMs (Touvron et al., 2023b; DeepSeek-AI, 2024). While previous practices often rely on human heuristics such as pre-defined rules and domain classification to filter data (Penedo et al., 2024b; Li et al., 2024), in this work we aim to directly identify data that could contribute effectively to learning various abilities.
@@
We are inspired by the belief that “compression represents intelligence” and the corresponding empirical finding recently (Huang et al., 2024), that the compression efficiency (i.e., the normalized loss) [^1] For a given model, the lossless compression efficiency, such as bits per character on certain data, is equal to the normalized loss on that data (Deletang et al., 2024; Huang et al., 2024). See §2.1 for further details. of a series of models on certain raw text, strongly correlates with their downstream benchmark performance, when the text data aligns with the benchmark domains. For example,  Huang et al. (2024) observed that model losses on raw GitHub code exhibit an almost linear correlation with performance on various coding tasks, whereas model losses on Common Crawl data are more indicative of performance on knowledge-intensive tasks. These insights suggest that compression efficiency on certain data is more reflective of model’s intelligence than that on other data. Motivated by this, we hypothesize that data on which compression more effectively represents intelligence also facilitates the learning of that respective intelligence more effectively.
@@
Building on this hypothesis, we first define the predictive strength of a given text to quantify how well the compression efficiency on it correlates with the models’ ability, then, we introduce predicative data selection (PreSelect), a data selection method based on the data’s predictive strength. Different from previous works that try to quantify the pretraining data influence requiring controlled training (Yu et al., 2024; Engstrom et al., 2024), our approach is extremely lightweight, and only utilizes open pretrained models in the wild without training any deep models in the process. After collecting a small seed dataset, a fastText (Joulin et al., 2017) classifier is trained to enable scalable data selection over a large corpus.
@@
While we are motivated independently by Huang et al. (2024), our approach shares similar intuition with Perplexity Correlation (Thrush et al., 2024) and can be viewed under the framework of Perplexity Correlation, which is the first to explore domain-level correlation-based methods in data selection. However, this work differs from Thrush et al. (2024) in several key aspects: (1) While they compute perplexity correlations for data domains (e.g., wikipedia.com), our method operates at a much finer granularity, directly operating on and selecting individual documents. [^2] We note that Thrush et al. (2024) preregistered document-level experiments as future tasks, which were not present as we released the first version of this paper. The updated version of Perplexity Correlation added document-level results concurrent to this work, but the final results are very different as we will dicuss in §2.3. As we will demonstrate in §3.4 and §4.3, this finer granularity leads to significant differences in the selected data and, consequently, in the final results; (2) Perplexity correlation suggested using a large number of (e.g., around 90) open-sourced models to approximate the correlation. However, we argue that using many open-sourced LLMs brings potential evaluation noises when dealing with different families of the models and causes computation inefficiency. Thus in our design, we employ 6 Llama models only during data selection. (3) They performed very limited experiments with a 160M-size model and compared with several baselines on limited benchmarks, while we train models with up to 3B parameters on up to 100B tokens and evaluate on 17 diverse benchmarks, a more realistic scale in a typical pretraining setting. The updated, concurrent version of Thrush et al. (2024) further added 1B-scale experiments assesed on comprehensive benchmarks, where the results indicate good performance on raw data pools but no improvement on high-quality, well-processes data pools – an observation distinct from our empirical results. PreSelect targets a well de-duplicated, high-quality pre-training corpus – which we believe represents a realistic setting – and shows the effectiveness of selecting data from high-quality data pools in §3.4.
@@ 2.2 Data’s Predictive Strength
We propose to select documents with high predictive strength scores, hypothesizing that the data that can be used to predict the model’s ability is also the data that contributes to learning it. Different from selection strategies based on pre-defined rules such as FineWeb-Edu (Penedo et al., 2024b) which prioritizes education-related documents, or DCLM (Li et al., 2024) which selects referring to supervised-finetuning data, our method is more principled and bypasses strong human heuristics.
@@ 2.3 The Overall Framework
Our core hypothesis is that the data on which the losses can help predict the performance well is the data that can contribute to training effectively, thus the documents’ predictive strength scores as defined in Eq. 1 are our metric to select data. However, computing Eq. 1 for every document is expensive since it requires computing losses of several models on all the data, thus we follow Li et al. (2024) and only select a small seed set of documents that are later used to train a fastText classifier (Joulin et al., 2017). As shown in Figure 2, we first randomly sample a small subset of data from the pre-training corpus, which is kept separate from the data used for later training, to compute the normalized losses. Specifically, we pick the 3,000 most frequent domains (e.g., wikipedia.org) in the corpus and randomly sample 300 examples for each domain for a wide coverage over the pre-training corpora, thus obtaining 900K samples in total. Then we choose the models from the Llama 1 and 2 series  (Touvron et al., 2023a, b) ranging from 7B to 65B parameters – in total 6 models are used to compute the predictive strength for the data. We do not choose various models from different families to compute the predictive strength on purpose – in our preliminary experiments, we found that models from various families introduce significant evaluation noises, for example, they are sensitive to prompts and the prompt that is good for one model may not be suitable for another. This issue is particularly salient given that we are evaluating base models that are generally more sensitive to prompts than instruction-tuned models. Such noises would dramatically affect the computation of predictive strength, and thus we found that data selected this way could not outperform a random data selection baseline in our initial trials. This design differs from Thrush et al. (2024) which employs diverse models to estimate perplexity correlation. Such distinction probably explains why perplexity correlation fails to improve random data selection baselines significantly on a high-quality, pre-filtered data pool, as shown in the concurrent version of their paper. We run inference of these models on the 900K documents to obtain the loss ranks of these models on each document respectively. Next, we obtain the downstream score ranks by directly using the averaged scores on a diverse set of 12 benchmarks from Huang et al. (2024).
@@ 3.1 Pre-training Corpus and Models
For a fair comparison and ease of processing, we follow and directly use the large data pool created in Li et al. (2024). Concretely, this data pool utilized a version of RefinedWeb that undergoes processing through resiliparse text extraction (Bevendorff et al., 2018), RefinedWeb’s heuristic filtering (Penedo et al., 2024d) and deduplication using bloom filters (Soldaini et al., 2024), solely filtered from Common Crawl. This large data pool has over 20 trillion tokens in total, which will serve as the source for us to sample smaller data pools to conduct data selection experiments in various scales. We refer to the pool as RefinedWeb in the following sections.
Following the suggested setting in Li et al. (2024), we randomly sample 80 billion, 300 billion, and 1 trillion tokens as the data selection pool for 400M, 1B, and 3B model training respectively and select 10% of the data for training unless otherwise specified, which corresponds to the Chinchilla optimal training data size (Hoffmann et al., 2024). For the models, we use a Llama architecture (Touvron et al., 2023a) unless otherwise specified. We also apply PreSelect on the C4 dataset to show the effectiveness of our method on different corpora and to compare with more data selection baselines.
@@ 3.2 Baselines
To make a fair comparison, we keep the data pool and training setting of baselines the same as PreSelect, such as model architecture and training hyper-parameters, while only the data selection part is different. We consider the following baselines: (1) Random, which randomly selects the documents; (2) PPL filtering, which keeps low-perplexity documents following CCNet (Wenzek et al., 2020); (3) FineWeb-Edu (Penedo et al., 2024b), which scores the education level of documents with an LLM and then train a small scorer to identify and select educational contents (Penedo et al., 2024b); (4) PPL Correlation (DD) (Thrush et al., 2024), which utilizes domain-level correlation estimation and performs domain-level selection. (5) PPL Correlation (DP) (Thrush et al., 2024), which utilizes domain-level correlation estimation and trains a fasttext based on domain-level positive/negative samples to perform page-level selection. As discussed before in §1, these two baselines are the most relevant to PreSelect while our approach does not rely on human heuristics to group documents and operates at the document level directly; and (6) DCLM (Li et al., 2024), which is the state-of-the-art pretraining data selection method that trains the fastText scorer using supervised fine-tuning (SFT) data as the positive data. For FineWeb-Edu and DCLM, we directly use their released BERT and fastText scorer to select documents in our pool. Due to the high cost of running all the baselines across all settings, we mainly include all of them in our setting of 1B model training for 30B tokens. In other experiment scales, we only compare to DCLM which is superior to other baselines. We provide more details about the model architecture and training hyper-parameters in Appendix B.3 and Appendix B.4 respectively.
@@ 3.4 Main Results
Compared with Perplexity Correlation (DD)which selects domains based on the correlation of each domain, it only achieves relatively marginal improvements over random selection and is significantly underperformed PreSelect by 4.8%. As we will further discuss in § 4.3, the removal of many domains leads to a significantly decreased diversity, thus, some tasks such as HellaSwag would even have a significant drop. In addition, we compare with Perplexity Correlation (DP) which perform page-level selection using fastText trained with domain-level positive/negative samples. We see that though it has 1.2% improvements over random baseline, it still underperformed PreSelect by 4.1%, showing the effectivness of using example-level correlation estimation. This difference is because by operating at the domain level, it may inadvertently include low-quality data from the high-correlation domains while overlooking high-quality data from unselected domains when building the fastText classifier.
@@ 4.2 FastText Feature Contribution
As shown in Table 3, for DCLM, the baseline that also uses a fastText classifier, the uni-gram features with high influence scores include (Why, 672), (Answer:, 620), (assistant., 567), (Question:, 448), (Describe, 368)… It is clearly observed that most positive uni-gram features have supervised fine-tuning data patterns and if the data in the pretraining corpus has such uni-gram, they have a high probability to be selected. While for PreSelect, the uni-gram features with high influence scores include (^ , 338), (MIT, 241), (Math, 194), (answer, 188), which widely covers many features appeared in code, education, and question answering domains. For example, “^” exists in many code snippets, “MIT” and “Math” exist in many education-related examples. These representative uni-gram features does not show in domain-level correlation estimation method, Perplexity Correlation (DP), where various diverse meaningless features are learned. We attribute this to the noise in the domain-level correlation calculation, where inadvertently include low-quality data from the high-correlation domains while overlooking high-quality data from unselected domains. The full learned positive and negative uni-grams alongside their influence score can be found in appendix A.3.2 which provides more insights into what word patterns are preferred by different methods.
@@ 5 Related Work
Given the growing available public sources such as 240 trillion tokens Common Crawl data released by (Li et al., 2024) which far exceeds the token budget feasible for training with current computational resources, data selection has emerged as a critical step to enhance data quality and optimize training efficiency nowadays. Existing approaches can broadly be categorized into rule-based methods and model-based methods.
5.1 Rule-based Data Selection
@@
5.2 Model-based Data Selection
In contrast to rule-based approaches, model-based data selection leverages trained models to score the data. For instance, CC Net retained the text that is assigned a low perplexity by a language model trained on Wikipedia (Wenzek et al., 2020), FineWeb-Edu trained a BERT classifier to prioritize educational contents (Penedo et al., 2024b), DsDm and MATES employed data influence models to assess the impact of specific data points on pretraining (Engstrom et al., 2024; Yu et al., 2024). And recently, DCLM (Li et al., 2024) also uses a fastText-based scorer which selects referring to supervised-finetuning data. These methods offer stronger performance than rule-based methods, however, they either require heavy computation or still contain strong human heuristics. Different from them, PreSelect is more principled and lightweight, thus bypassing strong human heuristics and offering adaptability. Furthermore, by operating at the document level, our method outperforms domain-level selection methods, such as those used by Thrush et al. (2024), in terms of granularity and effectiveness.

### v4

@@ Abstract
Language model pretraining involves training on extensive corpora, where data quality plays a pivotal role. In this work, we aim to directly estimate the contribution of data during pretraining and select pretraining data in an efficient manner. Specifically, we draw inspiration from recent findings showing that compression efficiency (i.e., normalized loss) of diverse models on certain text correlates strongly with their downstream performance, when the text domain aligns with the downstream benchmarks (Huang et al., 2024). Building on this observation, we hypothesize that data on which model losses are predictive of downstream abilities also contribute effectively to learning, which shares similar intuition with Thrush et al. (2024). To leverage this insight, we introduce predictive data selection (PreSelect), a lightweight and efficient data selection method that requires training and deploying only a fastText-based scorer. Through comprehensive experiments with 1B and 3B parameter models, we demonstrate that models trained on 30B tokens selected with PreSelect surpass the performance of the vanilla baseline trained on 300B tokens, achieving a 10x reduction in compute requirements. Furthermore, PreSelect significantly outperforms other competitive data selection baselines, such as DCLM and FineWeb-Edu on a scale of 3B models trained on 100B tokens. We open-source our trained data selection scorer along with the curated datasets at https://github.com/hkust-nlp/PreSelect.
@@ 1 Introduction
Large language model (LLM) pre-training typically requires training on a huge data source such as web crawl data where the existence of low-quality data leads to slow scaling law (Kaplan et al., 2020; Chowdhery et al., 2023). Given the growing available token budget and limited training budget nowadays, data selection has become a standard step in the pre-training stage. This step has played a critical role in enhancing data quality and optimizing training efficiency during the development of various LLMs (Touvron et al., 2023b; DeepSeek-AI, 2024). While previous practices often rely on human heuristics such as pre-defined rules and domain classification to filter data (Penedo et al., 2024b; Li et al., 2024a), in this work we aim to directly identify data that could contribute effectively to learning various abilities.
@@
We are inspired by the belief that “compression represents intelligence” and the corresponding empirical finding recently (Huang et al., 2024), that the compression efficiency (i.e., the normalized loss)For a given model, the lossless compression efficiency, such as bits per character on certain data, is equal to the normalized loss on that data (Deletang et al., 2024; Huang et al., 2024). See §2.1 for further details. of a series of models on certain raw text, strongly correlates with their downstream benchmark performance, when the text data aligns with the benchmark domains. For example,  Huang et al. (2024) observed that model losses on raw GitHub code exhibit an almost linear correlation with performance on various coding tasks, whereas model losses on Common Crawl data are more indicative of performance on knowledge-intensive tasks. These insights suggest that compression efficiency on certain data is more reflective of model’s intelligence than that on other data. Motivated by this, we hypothesize that data on which compression more effectively represents intelligence also facilitates the learning of that respective intelligence more effectively. A similar hypothesis “LLM losses on many pretraining texts are correlated with downstream benchmark performance” was proposed by Thrush et al. (2024) previously, but the empirical validation was limited and the effectiveness remained inconclusive.
@@
Building on this hypothesis, we first define the predictive strength of a given text to quantify how well the compression efficiency on it correlates with the models’ ability, then, we introduce predicative data selection (PreSelect), a data selection method based on the data’s predictive strength. Different from previous works that try to quantify the pretraining data influence requiring controlled training (Yu et al., 2024; Engstrom et al., 2024), our approach is extremely lightweight, and only utilizes open pretrained models in the wild without training any deep models in the process. After collecting a small seed dataset, a fastText (Joulin et al., 2017) is trained to enable scalable data selection.
@@
While we are motivated independently by Huang et al. (2024), our approach shares similar intuition with Perplexity Correlation (Thrush et al., 2024) and can be viewed under the framework of Perplexity Correlation, which is the first to explore domain-level correlation-based methods in data selection. However, this work differs from Thrush et al. (2024) in several key aspects: (1) While they compute perplexity correlations for data domains (e.g., wikipedia.com), our method operates at a much finer granularity, directly operating on and selecting individual documents.We note that Thrush et al. (2024) preregistered document-level experiments as future tasks, which were not present as we released the first version of this paper. The updated version of Perplexity Correlation added document-level results concurrent to this work, but the final results are very different as we will dicuss in §2.3. As we will demonstrate in §3.4 and §4.3, this finer granularity leads to significant differences in the selected data and, consequently, in the final results; (2) Perplexity correlation suggested using a large number of (e.g., around 90) open-sourced models to approximate the correlation. However, we argue that using many open-sourced LLMs brings potential evaluation noises when dealing with different families of the models and causes computation inefficiency. Thus in our design, we employ 6 Llama models only during data selection. (3) They performed very limited experiments with a 160M-size model and compared with several baselines on limited benchmarks, while we train models with up to 3B parameters on up to 100B tokens and evaluate on 17 diverse benchmarks, a more realistic scale in a typical pretraining setting. The updated, concurrent version of Thrush et al. (2024) further added 1B-scale experiments assesed on comprehensive benchmarks, where the results indicate good performance on raw data pools but no improvement on high-quality, well-processes data pools – an observation distinct from our empirical results. PreSelect targets a well de-duplicated, high-quality pre-training corpus – which we believe represents a realistic setting – and shows the effectiveness of selecting data from high-quality data pools in §3.4.
@@ 2.2 Data’s Predictive Strength
We propose to select documents with high predictive strength scores, hypothesizing that the data that can be used to predict the model’s ability is also the data that contributes to learning it. Different from selection strategies based on pre-defined rules such as FineWeb-Edu (Penedo et al., 2024b) which prioritizes education-related documents, or DCLM (Li et al., 2024a) which selects referring to supervised-finetuning data, our method is more principled and bypasses strong human heuristics.
@@ 2.3 The Overall Framework
Our core hypothesis is that the data on which the losses can help predict the performance well is the data that can contribute to training effectively, thus the documents’ predictive strength scores as defined in Eq. 1 are our metric to select data. However, computing Eq. 1 for every document is expensive since it requires computing losses of several models on all the data, thus we follow Li et al. (2024a) and only select a small seed set of documents that are later used to train a fastText classifier (Joulin et al., 2017). As shown in Figure 2, we first randomly sample a small subset of data from the pre-training corpus, which is kept separate from the data used for later training, to compute the normalized losses. Specifically, we pick the 3,000 most frequent domains (e.g., wikipedia.org) in the corpus and randomly sample 300 examples for each domain for a wide coverage over the pre-training corpora, thus obtaining 900K samples in total. Then we choose the models from the Llama 1 and 2 series  (Touvron et al., 2023a, b) ranging from 7B to 65B parameters – in total 6 models are used to compute the predictive strength for the data. We do not choose various models from different families to compute the predictive strength on purpose – in our preliminary experiments, we found that models from various families introduce significant evaluation noises, for example, they are sensitive to prompts and the prompt that is good for one model may not be suitable for another. This issue is particularly salient given that we are evaluating base models that are generally more sensitive to prompts than instruction-tuned models. Such noises would dramatically affect the computation of predictive strength, and thus we found that data selected this way could not outperform a random data selection baseline in our initial trials. This design differs from Thrush et al. (2024) which employs diverse models to estimate perplexity correlation. Such distinction probably explains why perplexity correlation fails to improve random data selection baselines significantly on a high-quality, pre-filtered data pool, as shown in the concurrent version of their paper. We run inference of these models on the 900K documents to obtain the loss ranks of these models on each document respectively. Next, we obtain the downstream score ranks by directly using the averaged scores on a diverse set of 12 benchmarks from Huang et al. (2024). We also explored how performance change when choosing one specific task (e.g. HellaSwag) as the target that elicits a different ranking to averaged score in Appendix A.7.2.
@@ 3.1 Pre-training Corpus and Models
For a fair comparison and ease of processing, we follow and directly use the large data pool created in Li et al. (2024a). Concretely, this data pool utilized a version of RefinedWeb that undergoes processing through resiliparse text extraction (Bevendorff et al., 2018), RefinedWeb’s heuristic filtering (Penedo et al., 2024d) and deduplication using bloom filters (Soldaini et al., 2024), solely filtered from Common Crawl. This large data pool has over 20 trillion tokens in total, which will serve as the source for us to sample smaller data pools to conduct data selection experiments in various scales. We refer to the pool as RefinedWeb in the following sections.
Following the suggested setting in Li et al. (2024a), we randomly sample 80 billion, 300 billion, and 1 trillion tokens as the data selection pool for 400M, 1B, and 3B model training respectively and select 10% of the data for training unless otherwise specified, which corresponds to the Chinchilla optimal training data size (Hoffmann et al., 2024). For the models, we use a Llama architecture (Touvron et al., 2023a) unless otherwise specified. We also apply PreSelect on the C4 dataset to show the effectiveness of our method on different corpora and to compare with more data selection baselines.
@@ 3.2 Baselines
To make a fair comparison, we keep the data pool and training setting of baselines the same as PreSelect, such as model architecture and training hyper-parameters, while only the data selection part is different. We consider the following baselines: (1) Random, which randomly selects the documents; (2) PPL filtering, which keeps low-perplexity documents following CCNet (Wenzek et al., 2020); (3) FineWeb-Edu (Penedo et al., 2024b), which scores the education level of documents with an LLM and then train a small scorer to identify and select educational contents (Penedo et al., 2024b); (4) PPL Correlation (DD) (Thrush et al., 2024), which utilizes domain-level correlation estimation and performs domain-level selection. (5) PPL Correlation (DP) (Thrush et al., 2024), which utilizes domain-level correlation estimation and trains a fasttext based on domain-level positive/negative samples to perform page-level selection. As discussed before in §1, these two baselines are the most relevant to PreSelect while our approach does not rely on human heuristics to group documents and operates at the document level directly; and (6) DCLM (Li et al., 2024a), which is the state-of-the-art pretraining data selection method that trains the fastText scorer using supervised fine-tuning (SFT) data as the positive data. For FineWeb-Edu and DCLM, we directly use their released BERT and fastText scorer to select documents in our pool. Due to the high cost of running all the baselines across all settings, we mainly include all of them in our setting of 1B model training for 30B tokens. (7) ScalingFilter (Li et al., 2024b), which uses the perplexity difference between a large model and a small model as the metric to select data, which we compared and discussed how intermediate level models actually help in Appendix A.7.1. In other experiment scales, we only compare to DCLM which is superior to other baselines. We provide more details about the model architecture and training hyper-parameters in Appendix B.3 and Appendix B.4 respectively.
@@ 3.4 Main Results
Compared with Perplexity Correlation (DD)which selects domains based on the correlation of each domain, it only achieves relatively marginal improvements over random selection and is significantly underperformed PreSelect by 4.8%. As we will further discuss in § 4.3, the removal of many domains leads to a significantly decreased diversity, thus, some tasks such as HellaSwag would even have a significant drop. In addition, we compare with Perplexity Correlation (DP) which perform page-level selection using fastText trained with domain-level positive/negative samples. We see that though it has 1.2% improvements over random baseline, it still underperformed PreSelect by 4.1%, showing the effectiveness of using example-level correlation estimation. This difference is because by operating at the domain level, it may inadvertently include low-quality data from the high-correlation domains while overlooking high-quality data from unselected domains when building the fastText classifier.
@@ 4.2 FastText Feature Contribution
As shown in Table 3, for DCLM, the baseline that also uses a fastText classifier, the uni-gram features with high influence scores include (Why, 672), (Answer:, 620), (assistant., 567), (Question:, 448), (Describe, 368)… It is clearly observed that most positive uni-gram features have supervised fine-tuning data patterns and if the data in the pretraining corpus has such uni-gram, they have a high probability to be selected. While for PreSelect, the uni-gram features with high influence scores include (ˆ , 338), (MIT, 241), (Math, 194), (answer, 188), which widely covers many features appeared in code, education, and question answering domains. For example, “ˆ” exists in many code snippets, “MIT” and “Math” exist in many education-related examples. These representative uni-gram features does not show in domain-level correlation estimation method, Perplexity Correlation (DP), where various diverse meaningless features are learned. We attribute this to the noise in the domain-level correlation calculation, where inadvertently include low-quality data from the high-correlation domains while overlooking high-quality data from unselected domains. The full learned positive and negative uni-grams alongside their influence score can be found in appendix A.3.2 which provides more insights into what word patterns are preferred by different methods.
@@ 5 Related Work
Rule-based Data Selection
@@
Model-based Data Selection
In contrast to rule-based approaches, model-based data selection leverages trained models to score the data. For instance, CC Net retained the text that is assigned a low perplexity by a language model trained on Wikipedia (Wenzek et al., 2020), Marion et al. (2023) and Ankner et al. (2024) also use perplexity as a practical and effective metric for data pruning, FineWeb-Edu trained a BERT classifier to prioritize educational contents (Penedo et al., 2024b), DsDm and MATES employed data influence models to assess the impact of specific data points on pretraining (Engstrom et al., 2024; Yu et al., 2024). And recently, DCLM (Li et al., 2024a) also uses a fastText-based scorer which selects referring to supervised-finetuning data. These methods offer stronger performance than rule-based methods, however, they either require heavy computation or still contain strong human heuristics. Different from them, PreSelect is more principled and lightweight, thus bypassing strong human heuristics and offering adaptability.